### v1

# Gluon Polarimetry with Energy-Energy Correlators

Abstract
We propose a novel method to probe gluon linear polarization via energy correlations in hard scattering processes. This approach exploits the characteristic $\cos 2\phi$ azimuthal modulation in single- and two-point energy correlations within jets initiated by polarized gluons. In contrast to conventional techniques that rely on $k_{t}$ resummation or intricate jet substructure observables, our method offers a theoretically robust and experimentally accessible avenue for gluon polarimetry. We perform an all-order analysis within the Ciafaloni-Catani-Fiorani-Marchesini (CCFM) formalism, incorporating coherent branching effects to achieve improved precision. Our predictions can be tested at current and future facilities, including the LHC, RHIC, HERA, and the EIC.
September 18, 2025
Introduction— Understanding the linear polarization of gluons inside unpolarized nucleons is essential for mapping their gluonic structure, a topic that has drawn significant interest from both small-$x$ and high-energy spin physics communities Mulders and Rodrigues (2001); Metz and Zhou (2011); Schafer and Zhou (2012); Boer et al. (2009, 2011); Qiu et al. (2011); Pisano et al. (2013); Boer et al. (2016, 2017); Kishore and Mukherjee (2019); Dominguez et al. (2012); Dumitru et al. (2015); Sun et al. (2011); Boer et al. (2012); Catani et al. (2014); Boer and den Dunnen (2014); Echevarria et al. (2015); Ma et al. (2013a); Gutierrez-Reyes et al. (2019); den Dunnen et al. (2014); Zhang (2014); Ma and Wang (2016); Altinoluk et al. (2024); Taels et al. (2022). The recent finding Bhattacharya et al. (2024), together with the observation that gluons exhibit 100% linear polarization at small $x$ in the dilute limit Metz and Zhou (2011), indicate that the helicity and orbital angular momentum (OAM) of gluons are maximally quantum entangled. This phenomenon has broad phenomenological implications, as evidenced by predicted substantial $\cos 2\phi$ azimuthal asymmetries in two particle production processes Boer et al. (2009, 2011); Qiu et al. (2011); Pisano et al. (2013); Boer et al. (2016); Schafer and Zhou (2012); Boer et al. (2017) and the transverse momentum distribution of Higgs boson at LHC Sun et al. (2011); Boer et al. (2012); Catani et al. (2014); Boer and den Dunnen (2014); Echevarria et al. (2015); Ma et al. (2013a); den Dunnen et al. (2014); Zhang (2014); Ma and Wang (2016); Gutierrez-Reyes et al. (2019). Despite these compelling signatures, directly measuring linear gluon polarization remains a major challenge. Traditional methods, such as those based on Transverse Momentum Dependent (TMD) factorization, suffer from two key limitations: the polarization signal is diluted by TMD evolution effects at low transverse momentum, and identical $\cos 2\phi$ asymmetries can also arise from final-state soft gluon emissions contaminating the extraction of Hatta et al. (2021a, b). To address these challenges, emerging observables like Energy-Energy Correlations (EECs) offer a novel and potentially more precise method for measuring linear gluon polarization.
Figure 1: The figure depicts a proton-proton collision in which a nearly collinear, linearly polarized gluon is emitted from an incoming parton involved in a hard scattering event. The magnified view on the right shows the resulting EEC within the jet cone initiated by the radiated gluon, clearly displaying the $\cos 2\phi$ anisotropy as revealed by our numerical calculations.
EEC based methods also leverage the $\cos 2\phi$ azimuthal asymmetry to probe linear gluon polarization. Specifically, a $\cos 2\phi$ azimuthal modulation in the squeezed limit of the three-point energy correlator within a jet Chen et al. (2021, 2022) provides a direct probe of the linear polarization of intermediate virtual gluons. Furthermore, the Nucleon EEC (NEEC) has emerged as a powerful tool for studying linearly polarized gluon distributions Li et al. (2023); Guo et al. (2025a). The scale dependence of the near-side EEC observables is governed by the DGLAP dynamics and is not sensitive to soft gluon radiations, yielding a theoretically cleaner analysis. Beyond jet substructure study, EEC, originally introduced in the 1970s Basham et al. (1978, 1979), have evolved into a versatile precision tool with applications ranging from hot nuclear matter Xing et al. (2025); Apolinário et al. (2025); Yang et al. (2024); Andres et al. (2023a, b, 2024a); Barata et al. (2024a); Barata and Szafron (2024); Andres et al. (2024b) to nucleon structure studies Liu and Zhu (2023); Guo et al. (2024, 2025a); Li et al. (2023); Liu et al. (2023); Bhattacharya et al. (2025); Mäntysaari et al. (2025); Kang et al. (2024a), and beyond Dixon et al. (2019); Ali et al. (2012); Tulipánt et al. (2017); Karapetyan (2019); Luo et al. (2019); Korchemsky (2020); Moult et al. (2020); Ebert et al. (2021); Gao et al. (2021); Li et al. (2021, 2022); Komiske et al. (2023); Neill et al. (2022); Craft et al. (2022); Chen et al. (2023); Devereaux et al. (2023); Yu et al. (2024); Schindler et al. (2023); Chen et al. (2024a); Kang et al. (2024a); Cao et al. (2024); Aglietti and Ferrera (2024); Singh and Vaidya (2024); Chen et al. (2024b); Shen et al. (2024); Liu et al. (2024); Kang et al. (2024b); Fu et al. (2024); Barata et al. (2024b); Andres et al. (2024c); Barata et al. (2024c); Alipour-fard and Waalewijn (2025); Acharya et al. (2024); Hayrapetyan et al. (2024); Zhao et al. (2025); Kang et al. (2025); Herrmann et al. (2025); Guo et al. (2025b); Chang et al. (2025); Lee and Stewart (2025); Chen et al. (2025); Gao et al. (2025)(for a nice review, see Ref. Moult and Zhu (2025) and references therein). Interestingly, Ref. Chen et al. (2024c) establishes a robust connection between NEEC and fracture functions.
Inspired by recent developments Chen et al. (2021, 2022), we introduce a novel approach to probe the linear polarization of collinear gluons emitted as initial-state radiation in $Z^{0}$ production process. Our method exploits the azimuthal dependence of the two-point energy-energy correlator measured within jets initiated by these gluons. The linear polarization, oriented along the gluon’s transverse momentum with respect to the nucleon beam, generates a characteristic $\cos 2\phi$ azimuthal asymmetry in EEC. Here, $\phi$ denotes the angle between the gluon’s transverse momentum ($P_{\perp}$) and the vector connecting the two detector pixels that contribute to the correlation. To achieve improved theoretical accuracy, we go beyond the conventional DGLAP formalism by incorporating coherent branching effects through the Ciafaloni-Catani-Fiorani-Marchesini (CCFM) framework Ciafaloni (1988); Catani et al. (1990a, b); Marchesini (1995). Notably, although partial cancellations occur between $g\rightarrow q\bar{q}$ and $g\rightarrow gg$ splittings, the resulting azimuthal asymmetry remains substantial. To enhance experimental accessibility, we further introduce a refined method to probe the linear polarization of gluon by employing the one-point energy correlation within the Winner-Takes-All (WTA) scheme, a well-established jet event shape observable Bertolini et al. (2014); Neill et al. (2017); Kang et al. (2017); Neill et al. (2019). The polarization signature is effectively encoded in a similar $\cos 2\phi$ modulation, where $\phi$ is defined as the angle between $P_{\perp}$ and the vector connecting the detector pixels to the WTA jet axis.
Our methodology shares the strengths of other EEC-based methods, notably in avoiding the contamination from final state soft gluon radiations and eliminating the complexities introduced by Sudakov resummation in such analyses. Compared to prior EEC based techniques, such as those relying on three point correlations Chen et al. (2021, 2022) or the NEEC probe Li et al. (2023), our method enhances precision and experimental feasibility by directly targeting the asymmetry within jets produced via polarized gluon fragmentation. With this robust framework, we aim to set a new standard for gluon polarimetry.
Formalism— Formally, the differential two-point EEC inside a gluon jet can be defined as Komiske et al. (2023),
$\frac{d\langle\text{EEC}\rangle}{d\theta\,d\phi}=\sum_{i,j}\delta(\theta_{ij}-\theta)\,\delta(\phi_{ij,g}-\phi)\,\frac{E_{i}E_{j}}{E_{g}^{2}},$ (1)
where the sum runs over all pairs of particles $(i,j)$ within the identified jet, and $E_{g}=\sum_{i}E_{i}$ is the total energy of the jet. Here, $\theta_{ij}$ is the angle between particles $i$ and $j$, and $\phi_{ij,g}$ denotes the azimuthal angle between the relative angle vector $\vec{\theta}_{ij}$ and the transverse momentum $\vec{P}_{\perp}$ of the jet.
To facilitate an all-order analysis, it is convenient to consider the $\theta$-integrated EEC, which can be expressed as a convolution of hard coefficients and the cumulant jet functions Dixon et al. (2019):
$\langle\text{EEC}\rangle=$ $\sum_{i=q,g}\int dx\,x^{2}H_{i}(x,\mu)\,J_{i}(\ln x^{2}\kappa)$ (2)
where $\kappa=\theta^{2}E_{g}^{2}/\mu^{2}$, and $x$ is the momentum fraction of the jet carried by the parton $i$. $J_{i}$ denotes the standard (unpolarized) cumulant jet function Dixon et al. (2019). $H_{i}$ is the hard coefficient of resolving a parton carrying momentum fraction $x$ within the gluon jet at the scale $\mu$, which will be specified later.
As discussed above, there is a direct correlation between the gluon branching plane and its polarization vector. In the context of QCD evolution, this process can be naturally divided into two distinct stages: First, as the parton shower develops, the linear polarization of gluons is progressively reduced due to scale evolution effects. Second, when a linearly polarized gluon with offshellness $E_{g}\theta$ undergoes a single splitting, the resulting energy distribution exhibits a characteristic azimuthal dependence. Taking these effects into account, we arrive at a factorized expression for the $\cos 2\phi$-weighted EEC:
$\int d\phi\,\frac{d\langle\text{EEC}\rangle}{d\theta\,d\phi}\cos 2\phi=\int dx\,x^{2}H_{T}(x,\mu)$ (3)
$\times\frac{\alpha_{s}}{2\pi}\frac{1}{\theta}\int\!dyy(1-y)\left[P_{gg}^{2\phi}(y)+2n_{f}P_{gq}^{2\phi}(y)\right]J_{g,T}(\ln x^{2}\kappa)$
Here, $J_{g,T}$ is the polarization-dependent jet function, whose operator definition can be found in Ref. Chen et al. (2021). The relevant azimuthal-dependent splitting functions are:
$P_{gg}^{2\phi}(y)=2C_{A}\,y(1-y),\quad P_{qg}^{2\phi}(y)=-y(1-y)$ (4)
$H_{T}$ is the corresponding hard coefficient for producing a linearly polarized gluon.
To proceed further, we note that the observable studied here is closely related to the NEEC Liu and Zhu (2023); Li et al. (2023); Guo et al. (2025a), since we focus on gluon radiation in the forward region along the beam direction, where $P_{\perp}\ll Q$. We therefore adopt a similar factorization approach, in which the initial-state collinear gluon radiation is separated from the subsequent hard scattering process. As an example, consider $Z^{0}$ boson production, where an incoming quark emits a gluon before the hard interaction:
$\frac{d\sigma_{U}}{d\xi dy_{Z}d^{2}P_{\perp}}$ $=\sigma_{0}\,\frac{\alpha_{s}}{2\pi}\,C_{F}\,\frac{1}{P_{\perp}^{2}}\,\frac{1+(1-\xi)^{2}}{\xi},$ (5)
$\frac{d\sigma_{T}}{d\xi dy_{Z}d^{2}P_{\perp}}$ $=\sigma_{0}\,\frac{\alpha_{s}}{2\pi}\,C_{F}\,\frac{1}{P_{\perp}^{2}}\,\frac{2(1-\xi)}{\xi},$ (6)
where $\sigma_{0}$ is the Born-level cross section for $Z^{0}$ production, and $\xi$ denotes the longitudinal momentum fraction carried by the gluon jet. The ratio $\sigma_{T}/\sigma_{U}$ quantifies the degree of linear polarization of the emitted gluon. Therefore, a natural choice for hard coefficients reads $H_{i}(x,\mu=\theta_{0}E_{g})=\delta_{ig}\delta(x-1)$ and $H_{T}(x,\mu=\theta_{0}E_{g})=\sigma_{T}/\sigma_{U}\delta(x-1)$, with $\theta_{0}\sim{\cal O}(1)$.
In the DGLAP formalism, the evolution equations of the jet functions are given by
$\frac{\partial J_{i}(\ln\kappa)}{\partial\ln\mu^{2}}=\frac{\alpha_{s}}{2\pi}\sum_{j}\int_{0}^{1}dyy^{2}P_{ij}(y)J_{j}(\ln(y^{2}\kappa)),$ (7)
$\frac{\partial J_{g,T}(\ln\kappa)}{\partial\ln\mu^{2}}=\frac{\alpha_{s}}{2\pi}\int_{0}^{1}dyy^{2}P_{gg}^{T}(y)J_{g,T}(\ln(y^{2}\kappa)).$ (8)
Here $i,j$ denote the parton flavors, and $P_{ij}$ is the unpolarized splitting function. $P_{gg}^{T}(y)$ is the linearly polarized gluon splitting function, which at the LO reads Sather and Schmidt (1990); Ma et al. (2013b)
$P^{T}_{gg}(y)=2C_{A}\frac{y}{(1-y)_{+}}+\beta_{0}\delta(1-y),$ (9)
with $\beta_{0}=(33-2n_{f})/6$.
Having discussed two-point EECs, we now consider the single-point energy correlation, or jet event shape. This observable quantifies the energy distribution within a jet by measuring the correlation between the energy of a radiated parton and the jet core itself. A key advance in this context is the Winner-Take-All (WTA) jet axis reconstruction scheme Bertolini et al. (2014); Neill et al. (2017); Kang et al. (2017); Neill et al. (2019). Unlike conventional algorithms that define the jet axis as the sum of all constituent momenta, the WTA algorithm assigns the jet axis to the direction of the hardest particle at each recombination step. This procedure effectively eliminates recoil effects from soft gluon radiation, resulting in a recoil-free jet axis that enables more precise measurements of jet substructure and polarization-sensitive observables.
Within the WTA scheme, the unpolarized and linearly polarized single-point jet functions are denoted by $D_{g/q}$ and $D_{g,T}$, respectively. In the DGLAP formalism, their scale evolution is governed by
$\frac{\partial D_{i}(\ln\kappa)}{\partial\ln\mu^{2}}=\frac{\alpha_{s}}{2\pi}\sum_{j}\int_{\frac{1}{2}}^{1}dyyP_{ij}(y)D_{j}(\ln y^{2}\kappa),$ (10)
$\frac{\partial D_{g,T}(\ln\kappa)}{\partial\ln\mu^{2}}=\frac{\alpha_{s}}{2\pi}\int_{\frac{1}{2}}^{1}dyyP_{gg}^{T}(y)D_{g,T}(\ln y^{2}\kappa).$ (11)
Here, the lower integration limit $y=1/2$ arises from the kinematic constraint in the WTA scheme. Additionally, note that the evolution equation involves a single power of $y$ (i.e., $y$), in contrast to the two-point EEC case where $y^{2}$ appears. Note that the one point energy correlation under investigation differs from that studied in Ref. Mi and Wang (2025).
Although the DGLAP formalism is the standard approach for describing the scale evolution of jet functions, it does not account for quantum coherence effects among gluon emissions. As a result, it tends to overcount soft gluons emitted at large angles and fails to capture the angular correlations that are crucial for an accurate description of parton showers and jet substructure. The CCFM formalism Ciafaloni (1988); Catani et al. (1990a, b); Marchesini (1995) overcomes these limitations by implementing angular ordering: each successive emission in the final-state parton shower must occur at a smaller opening angle than the previous one. This approach not only regulates infrared divergences more effectively—ensuring, for instance, that the EEC remains finite as the angular separation $\theta\to 0$—but also yields a more realistic description of energy flow and correlations within jets.
In the original CCFM approach, it is necessary to introduce a parton distribution that depends on the transverse momentum in order to resum large logarithms at small $x$. By analogy, we also define a transverse momentum dependent jet function, $J_{g}(\ln\kappa,k_{t})$, which satisfies the normalization condition $\int d^{2}k_{t}\,J_{g}(\ln\kappa,k_{t})=J_{g}(\ln\kappa)$. This construction allows us to directly apply the CCFM formalism to the evolution of the jet functions.
Following Refs. Marchesini (1995); Jung (2002); Jung et al. (2010), and focusing for simplicity on the gluon-to-gluon channel, the CCFM evolution equation for the unpolarized gluon jet function can be written in differential form as:
$\frac{\partial}{\partial\ln\mu^{2}}\frac{J_{g}(\ln\kappa,k_{t})}{\Delta_{s}(\mu^{2})}=\frac{\alpha_{s}}{2\pi}\int_{\Lambda/\mu}^{1-\Lambda/\mu}dy\,y^{2}$ (12)
$\qquad\times\int_{0}^{2\pi}\frac{d\phi^{\prime}}{2\pi}\frac{\tilde{P}_{gg}(y,\mu^{2},k_{t})}{\Delta_{s}(\mu^{2})}J_{g}(\ln\kappa,|\vec{k}_{t}+(1-y)\vec{\mu}|),$
where $\mu^{2}=\vec{\mu}^{2}$ sets the maximum allowed angle for gluon emission, $\theta_{\text{max}}=\mu/E_{g}$, and $\Lambda$ is an infrared cutoff. The angle $\phi^{\prime}$ denotes the azimuthal separation between $\vec{k}_{t}$ and $\vec{\mu}$. The Sudakov form factor $\Delta_{s}$ resums the effects of virtual and unresolved real emissions. Importantly, due to the implementation of angular ordering in CCFM, the $y$-dependence in the logarithm is eliminated.
In the CCFM framework, the splitting function $\tilde{P}_{gg}$ is unregularized, retaining only the singular terms $1/y$ and 1/(1y)1/(1-y). In the original context, a non-Sudakov form factor $\Delta_{ns}$ is introduced to resum logarithms of the form $\ln(1/y)$. However, for energy-energy correlation (EEC) observables, the energy-weighting factors $y$ or $y^{2}$ suppress the $1/y$ divergence, making small-$x$ resummation via the non-Sudakov factor unnecessary. Consequently, the unregularized splitting function for the unpolarized case simplifies to $\tilde{P}_{gg}(y)=2C_{A}\left[\frac{y}{1-y}+\frac{1-y}{y}+y(1-y)\right],$ which is independent of the transverse momentum $k_{t}$.
With this simplification, we can integrate over $k_{t}$ to obtain an evolution equation for the inclusive jet function:
$\frac{\partial}{\partial\ln\mu^{2}}\frac{J_{g}(\ln\kappa)}{\Delta_{s}(\mu^{2})}=\frac{\alpha_{s}}{2\pi}\int_{\Lambda/\mu}^{1-\frac{\Lambda}{\mu}}\!\!\!dyy^{2}\tilde{P}_{gg}(y)\frac{J_{g}(\ln\kappa)}{\Delta_{s}(\mu^{2})}.$ (13)
The corresponding Sudakov form factor is given by
$\Delta_{s}(\mu^{2})=$ (14)
$\exp\Bigg\{\!-\!\int_{4\Lambda^{2}}^{\mu^{2}}\!\!\frac{d\mu^{\prime 2}}{\mu^{\prime 2}}\frac{\alpha_{s}}{4\pi}\int_{\Lambda/\mu^{\prime}}^{1-\frac{\Lambda}{\mu^{\prime}}}\!\!dy[\tilde{P}_{gg}(y)+2n_{f}\tilde{P}_{qg}(y)]\Bigg\}.$
where $\tilde{P}_{qg}(y)$ is the unregularized splitting function for the gluon-to-quark channel. This form closely parallels the Sudakov factor used in parton shower algorithms.
Extending the above discussion to the case of linearly polarized gluons, we find that the evolution of the corresponding jet function $J_{g,T}$ within the CCFM framework takes a similar form to the unpolarized case,
$\frac{\partial}{\partial\ln\mu^{2}}\frac{J_{g,T}(\ln\kappa)}{\Delta_{s}(\mu^{2})}=\frac{\alpha_{s}}{2\pi}\int_{\Lambda/\mu}^{1-\frac{\Lambda}{\mu}}\!\!\!dy\,y^{2}\frac{\tilde{P}_{gg}^{T}(y)}{\Delta_{s}(\mu^{2})}J_{g,T}(\ln\kappa),$ (15)
where the unregularized splitting function for linearly polarized gluons is given by $\tilde{P}_{gg}^{T}(y)=2C_{A}\frac{y}{1-y}$.
The CCFM formalism can be directly extended to single-point energy correlations in the WTA scheme. In the next section, we examine the phenomenological consequences of these improvements, particularly for azimuthal asymmetries from linearly polarized gluons.
Phenomenological Analysis— We now turn to the phenomenological analysis of EECs and their sensitivity to gluon polarization.
The boundary conditions for the jet functions are set at a reference angle $\theta_{0}$ (e.g., $\theta_{0}=0.3$), corresponding to a factorization scale $\mu_{0}=P_{\perp}=30$ GeV for gluon energy $E_{g}=90$ GeV, as motivated by the associated $Z^{0}$ production process. Specifically, we impose
$J_{g}(0)=1,\ J_{q}(0)=0,\ J_{g,T}(0)=1,$ (16)
The quark jet function is initialized to zero, reflecting the fact that quark contributions arise only from subsequent branchings.
Collecting all pieces together, the azimuthal asymmetry for the EEC, which encodes the effect of gluon linear polarization, is given by $\langle\cos 2\phi\rangle=A(\theta)H_{T}/H$ where the analyzing power $A(\theta)$ is defined as
$A(\theta)\equiv\frac{\int\!dyy(1-y)\left[P_{gg}^{2\phi}+2n_{f}P_{qg}^{2\phi}\right]\!J_{g,T}}{\int\!dyy(1-y)\!\left\{\left[\tilde{P}_{gg}+2n_{f}\tilde{P}_{qg}\right]\!J_{g}+\!\left[\tilde{P}_{qq}+\tilde{P}_{gq}\right]\!J_{q}\right\}},$ (17)
with the limits of the $y$ integration being set to $\int_{\Lambda/\mu_{0}}^{1-\Lambda/\mu_{0}}$ in order to account for the coherent branching effects. For the single-point energy correlation in the WTA scheme, the expression of the asymmetry remains analogous. The analyzing power $A(\theta)$ serves as a key measure of the observable’s sensitivity to gluon polarization, and is central to our phenomenological study.
We now present numerical results for the unpolarized distribution and the $\cos 2\phi$ azimuthal modulations. Figure 2 shows the self-normalized unpolarized angular distributions for the EEC: $Z_{g}(\theta)=\partial J_{g}(\ln\kappa)/\partial\theta$ in the region $\theta\in[0.01,0.3]$, using the boundary conditions specified above. We compare results obtained from standard DGLAP evolution with those from the CCFM formalism, using infrared cutoffs $\Lambda=0.3$ GeV and $\Lambda=0.7$ GeV. Notably, our calculations qualitatively reproduce the characteristic plateau behavior of EEC. Therefore, the infrared cutoff inherent in the CCFM formalism provides additional flexibility to describe the nonperturbative behavior of the near-side EEC.
Figure 2: Self-normalized unpolarized angular distributions for the EEC in the region $\theta\in[0.01,0.3]$, calculated with the boundary conditions $J_{g}(0)=1$ and $J_{q}(0)=0$ at $\mu=30$ GeV. Results are shown for both standard DGLAP evolution and the CCFM formalism with infrared cutoffs $\Lambda=0.3$ GeV and $\Lambda=0.7$ GeV.
Figure 3 illustrates the analyzing power $A(\theta)$, which quantifies the maximal $\cos 2\phi$ azimuthal asymmetry, for both single-point and two-point energy-energy correlations (EECs) as a function of the angle $\theta$ in the range $0.01$ to $0.3$. The shaded error bands indicate the impact of varying the infrared cutoff $\Lambda$ between $0.3$ GeV and $0.7$ GeV. The results demonstrate that the azimuthal asymmetry is sizable and should be accessible to measurement at the LHC.
Figure 3: Gluon polarization analyzing power $A(\theta)$, representing the maximal azimuthal asymmetry $\langle\cos{2\phi}\rangle$, for single-point and two-point energy correlations with $n_{f}=5$, plotted as a function of $\theta$ in the range $0.01$ to $0.3$ at $\mu=P_{\perp}=30$ GeV. The error bands show the effect of varying the infrared cutoff $\Lambda$ from $0.3$ GeV to $0.7$ GeV.
Summary— We introduce a novel method to probe the linear polarization of gluons by analyzing the energy distribution within jets produced by initial-state radiation in Drell-Yan-like processes. Gluon polarization appears as a characteristic $\cos 2\phi$ modulation in the azimuthal dependence of the EEC and jet event shapes within the WTA scheme.
Our approach is innovative both phenomenologically, as it provides direct sensitivity to gluon polarization without relying on delicate jet substructure analysis, and theoretically, through the use of the CCFM formalism to incorporate coherent branching effects. Notably, the CCFM formalism naturally captures the plateau behavior observed in the EEC, providing a unified description of both perturbative and nonperturbative effects. While our study centers on Drell-Yan processes, the technique extends seamlessly to other hard scattering processes, such as DIS and mid-rapidity di-jet production accompanied by initial-state gluon radiation as shown in Fig. 4. The predicted asymmetries can be validated with existing data from HERA, RHIC, and the LHC, with further tests anticipated at the future EIC. In summary, this approach establishes a new framework for gluon polarimetry with EECs in the perturbative large-angle regime, while shedding new light on the transition from non-perturbative to perturbative collinear dynamics in the polarization-dependent case.
Figure 4: Schematic illustration of linearly polarized gluons from initial-state radiation in DIS (left) and di-jet production (right), showing characteristic energy patterns in jet cones that enable probing gluon polarization.
Acknowledgements: We thank Xiao-hui Liu, Ting Lin and Jin-long Zhang for helpful discussions. The work of J. Z. is supported in part by the National Science Foundations of China under Grant No. 12175118, and No. 12321005. This work is also in part supported by Natural Science Foundation of China under grant No. 12405156 and No. 11505080, the Shandong Province Natural Science Foundation under grant No. 2023HWYQ-011, No. ZFJH202303 and No. ZR2018JL006, and the Taishan fellowship of Shandong Province for junior scientists.